OpenStudy (anonymous):

HELP!!!! In biology, the function defined by g(x) = (5x)/(2+x) gives the growth rate of a population for x units of available food. a) What amount of food (in appropriate units) would produce a growth rate of 3 units of growth per unit of food? b) What is the growth rate if no food is available?

OpenStudy (lilg132):

5 * 3 = 15 2 + 3 = 5 g(3) = 15 / 5 g(3) = 3

OpenStudy (anonymous):

i dont like biology, sorry :D

OpenStudy (anonymous):

calculate g(0) for no food

OpenStudy (lilg132):

g(0) = 5(0) / (2 + 0) g(0) = 2

OpenStudy (anonymous):

yay! thanks lilg132, but the answer for b is 0 though

OpenStudy (anonymous):

yeah 0/2=0 :D

OpenStudy (lilg132):

ooooops !my bad :p